OpenStudy (anonymous):

Use multiplication, division, addition and subtraction and at least one set of parenthesis to write an expression that simplifies 7. Please explain step by step and explain each step as you simplify the expression.

OpenStudy (anonymous):

what does "simplifies 7" mean?

OpenStudy (anonymous):

you have to write an expression that gives you the answer of seven

OpenStudy (mr.math):

I think he/she means that the expression can be simplified to 7.

OpenStudy (anonymous):

ooooooooooh like \[4+3\] or \[\frac{1}{2}(8+6)\]

OpenStudy (anonymous):

since when did "equals" become "simplify" ??? bad math teacherese

OpenStudy (anonymous):

start with 7 and work backwards \[7=10-3=10-(5-2)\] etc until you have included all operations you need

OpenStudy (anonymous):

English is a crazy language! Simplify and equals can mean the same thing, just like pail or bucket ....